# Recent Advances in Online Hate Speech Moderation: Multimodality and the Role of Large Models

### v2

@@ Abstract
In the evolving landscape of online communication, moderating hate speech (HS) presents an intricate challenge, compounded by the multimodal nature of digital content. This comprehensive survey delves into the recent strides in HS moderation, spotlighting the burgeoning role of large language models (LLMs) and large multimodal models (LMMs). Our exploration begins with a thorough analysis of current literature, revealing the nuanced interplay between textual, visual, and auditory elements in propagating HS. We uncover a notable trend towards integrating these modalities, primarily due to the complexity and subtlety with which HS is disseminated. A significant emphasis is placed on the advances facilitated by LLMs and LMMs, which have begun to redefine the boundaries of detection and moderation capabilities. We identify existing gaps in research, particularly in the context of underrepresented languages and cultures, and the need for solutions to handle low-resource settings. The survey concludes with a forward-looking perspective, outlining potential avenues for future research, including exploring novel AI methodologies, the ethical governance of AI in moderation, and developing more nuanced, context-aware systems. This comprehensive overview aims to catalyze further research and foster a collaborative effort towards more sophisticated, responsible, and human-centric approaches to HS moderation in the digital era.111 WARNING: This paper contains offensive examples.
**footnotetext: Both authors contributed equally to this work
@@ 1 Introduction
In the era of rapid information exchange and digital connectivity, the rise of hate speech (HS) presents a significant challenge with profound implications for global societies. Hate speech, which is any communication demeaning a person or group based on social or ethnic characteristics, undermines social harmony and individual safety, both online and offline Awan and Zempi (2016); Lupu et al. (2023); Wiedlitzka et al. (2023). The recent Israel-Hamas conflict has notably escalated both anti-Muslim and anti-Semitic sentiments worldwide, evidenced by the trending of hashtags such as #HitlerWasRight and #DeathToMuslim on social media platform X. [^2] https://www.nytimes.com/2023/11/15/technology/hate-speech-israel-gaza-internet.html Additionally, The Council on American-Islamic Relations reported receiving 774 help requests and bias reports from Muslims in the USA within a 16-day period. [^3] https://www.cair.com/press_releases/cair-reports-sharp-increase-in-complaints-reported-bias-incidents-since-107/ While digital interconnectivity facilitates swift information sharing, it simultaneously amplifies the spread and impact of HS, transcending geographical boundaries.
@@
Technological advancements have transformed the expression of HS, leading to its manifestation in various novel forms. Traditionally, HS was predominantly text-based, found in written materials Ana (1999), or verbalized in posts, broadcasts, and public speeches Nielsen (2002). However, the digital era has ushered in more complex and subtle variants of HS, engaging multiple sensory modalities. A notable instance is vision-language HS, which fuses visual elements with textual content, commonly disseminated through captioned images and memes Kiela et al. (2020); Fersini et al. (2022); Bhandari et al. (2023). Video-based HS, another emerging form, amalgamates text, visuals, and audio, creating a multi-faceted and potentially more influential mode of communication Das et al. (2023). Figure 1 exemplifies various HS forms targeting immigrants, underscoring animosity towards individuals of diverse nationalities. The text-based approach overtly projects hostile attitudes towards immigrants in the host country. In vision-language HS, visual (e.g., a person preparing to shoot) and textual elements (e.g., sighting an illegal immigrant mowing the lawn) jointly convey antagonism. The figure also includes a music parody, integrating derogatory visuals with discriminatory audio lyrics, to showcase contempt for immigrants.
While the digital landscape has seen increasing diversity in HS forms, existing research surveys Rini et al. (2020); Chhabra and Vishwakarma (2023); Subramanian et al. (2023) predominantly concentrate on textual HS, often overlooking the complexity introduced by multimodal content. This survey aims to address this critical gap. Unlike prior surveys, we provide a comprehensive analysis of HS across various digital platforms, encompassing not only text but also visual, auditory, and combined multimodal expressions. We delve into the nuanced ways in which HS manifests in these different formats, offering insights into their unique characteristics and the challenges they pose for moderation.
Furthermore, our survey underscores the pivotal role of large language models (LLMs) and large multimodal models (LMMs) in moderating HS. These advanced models, which can process and interpret multiple types of data simultaneously, are crucial in understanding and countering the multifaceted nature of modern HS. We critically assess the existing solutions, evaluating their effectiveness and highlighting areas for improvement. Our survey presents a novel perspective on HS moderation, shifting the focus from traditional text-based analyses to multimodal approaches.
For this survey, we methodically examined research on moderating various types of hate speech, including text, images, videos, and audio. We searched with keywords like ‘hate speech’, ‘multimodal hate speech’, ‘harmful memes’, etc., on various scholarly platforms such as Google Scholar, DBLP, IEEE Xplore, and ACM Digital Library.
@@ 2 Hate Speech
Figure 2: Typology of HS based on modalities and tasks. The dark blue boxes are mature research areas with multiple studies; light grey boxes are ongoing research areas, while the hat can be found, and hatched boxes are unexplored topics.
HS takes various forms — written text, images, spoken words, and multimedia content — each posing risks of violence, animosity, or prejudice against specific groups. This section reviews existing literature on HS, categorizing it into text-based, image-based, video-based, and audio-based types. Figure 2 illustrates the range of online HS forms. Additionally, Table 1 lists publicly accessible HS datasets for different modalities, providing researchers with essential resources for data collection and analysis, and supporting comprehensive and consistent research in HS detection and moderation.
@@ 2.1 Text-based Hate Speech
Text-based HS encompasses written or typed expressions manifesting across various digital formats, including social media posts Waseem and Hovy (2016); Founta et al. (2018). Researchers addressing online HS have compiled datasets to detect and analyze its complexity and multifaceted nature, as detailed in existing literature. These studies explore diverse aspects of hateful and derogatory language, focusing on themes such as implicit HS ElSherief et al. (2021), targeted groups Kennedy et al. (2018), and types of attacks ElSherief et al. (2021). As detection models evolve, it becomes imperative to understand and elucidate their decision-making processes, leading to the development of datasets for implicit statements and methodologies for bias analysis Sap et al. (2020); ElSherief et al. (2021). Additionally, some research shifted towards proactive strategies, including countering HS Masud et al. (2022).
The detection of text-based HS is fraught with significant challenges. A primary obstacle is contextual understanding; HS often subtly or covertly conveys its message through sarcasm, irony, or cultural references. Linguistic variations, including slang, dialects, and unconventional language forms, exacerbate this intricacy, posing difficulties for automated recognition systems. Additionally, the multilingual nature of online content necessitates language-specific tools or models for effective detection across diverse linguistic contexts. A further complication is the rapid evolution of online language, which can quickly render existing detection models outdated.
@@ 2.2 Image-based Hate Speech
Image-based HS utilizes visual elements, such as photographs, cartoons, and illustrations, to propagate hate or discrimination against specific groups. A common manifestation of this type is memes, which typically consist of images combined with short overlaid text. Although memes often serve humorous or satirical purposes, they are increasingly used to spread hateful content online Kiela et al. (2020). The viral nature of memes on digital platforms further amplifies their impact, allowing them to reach vast audiences rapidly. To address the proliferation of hateful memes, recent studies have focused on developing datasets for their detection Kiela et al. (2020); Gomez et al. (2020) and for identifying specific targets and types of attacks within these memes Fersini et al. (2022). Beyond detection, new approaches are being pursued to analyze and mitigate bias in image-based HS detection models Hee et al. (2023); Lin et al. (2023). Additionally, new methodologies are emerging to counteract HS transmitted through memes Van and Wu (2023).
Detecting image-based HS presents challenges due to the nuanced expression of offensive messages in visual content. Images, often embedding symbols, memes, or culturally specific visual cues, require deep cultural and contextual understanding for accurate interpretation. The combination of visual elements and text can subtly imply meanings not immediately evident Kiela et al. (2020). For example, Figure 1 depicts a man with a gun and text suggesting hostility towards immigrants. Differentiating humor from hate in memes is particularly challenging, influenced by varying cultural, societal, and personal perspectives Schmid (2023).
@@ 2.3 Video-based Hate Speech
Video-based HS presents a complex challenge, comprising a blend of visuals, audio tracks, and/or textual elements. This form of HS ranges from professionally produced propaganda to amateur videos on social media platforms like YouTube and TikTok. The engaging nature of video content and its easy dissemination across digital networks significantly heighten its potential for harm. Echoing the concerns of image-based HS, video-based HS also contributes to the normalization of hateful ideologies and can profoundly influence public opinion. Contemporary research in this domain primarily focuses on identifying video-based HS and categorizing its various subtypes Wu and Bhandary (2020). Recent studies have explored detecting hateful segments within videos alongside multi-dimensional analysis. Nonetheless, the body of research on video-based HS is less developed than text-based and image-based HS, particularly in areas such as analyzing and mitigating model bias, elucidating decision-making processes, and devising counterstrategies. These gaps, likely stemming from the rapid pace of technological advancements and evolving digital trends, underscore the need for further research to promote a more harmonious online environment.
Identifying HS within videos is a demanding and resource-intensive endeavor. This complexity stems from the amalgamation of various elements, such as text, images, and audio, within a single medium. Each component can independently harbor HS, thereby compounding the detection process. The duration of videos further exacerbates this challenge, as longer content necessitates more extensive review and analysis, with potential shifts in context over time. Moreover, subtle visual cues and sophisticated editing techniques can be employed to discreetly embed hate messages, making their detection by automated tools particularly challenging. Additionally, the analysis of video content requires considerable computational resources and time, posing a substantial challenge for organizations tasked with detecting and addressing HS in video formats.
@@ 2.4 Audio-based Hate Speech
Audio-based HS entails the analysis of sound waves to discern elements such as pitch, intonation, and the contextual meaning of spoken words. This form of HS can originate from a variety of audio channels, including real-time conversations, podcasts, broadcasts, and other forms of audio media. The methodologies for addressing audio-based HS are diverse, targeting different facets of the issue. For instance, Barakat et al. Barakat et al. (2012) employed a straightforward keyword-based approach to identify segments of HS, while Wazir et al. Wazir et al. (2020) engaged in a detailed classification of offensive categories in audio-based HS, showcasing a nuanced method of understanding and categorizing this form of HS. This research area is still in its developmental stages, partly due to the scarcity of dataset. Nonetheless, recognizing the variety and significance of the approaches and techniques employed in this field is imperative. This recognition not only sheds light on the current state of research but also illuminates potential avenues for future exploration.
@@ 3 Methodology
Mod. Dataset Task Labels Source # Records
Text WZ-LS Waseem and Hovy (2016) Det.
[M.C.] Sexism, Racism, Neither
Twitter 16,914
GHC Kennedy et al. (2018) Det.
[M.C.] VO, HD, CV [B] Implicit, Explicit [M.C.] Hate Targets
Forums 27,665
Stormfront de Gibert et al. (2018) Det.
[B] Hateful
StormFront 9,916
DT Davidson et al. (2017) Det.
[M.C.] Hateful, Offensive, Neither
Twitter 24,802
HatEval Basile et al. (2019) Det.
[B] Hateful [B] Target [B] Aggressive
Twitter 19,600
Founta Founta et al. (2018) Det.
[M.C.] Offensive, Abusive, Hateful Speech, Aggressive, Cyberbullying, Spam, Normal
Twitter 80,000
SBIC Sap et al. (2020) Det.
[B] Offensive [B] Intent [B] Lewd [B] Group [B] Hate Targets [B] In-Group
Mixed 44,671
Expl.
[F.T.] Implied Statement
IHC ElSherief et al. (2021) Det.
[M.C.] Implicit, Explicit, Non-Hate [M.C.] Grievance, Incitement, Inferiority, Irony, Stereotypical, Threatening, Others
Twitter 22,584
Expl.
[F.T.] Implied Statement
HateXplain Mathew et al. (2021) Det.
[M.C] Hate, Offensive, Normal [M.C.] Hate Targets
Mixed 20,148
Expl.
[M.L] Text Rationales/Snippets
DynaHate Vidgen et al. (2021) Det.
[B] Hateful, [M.C] Animosity, Derogation, Dehumanization, Threatening, Support, [M.C.] Hate Targets
H-M Adv 41,134
ToxiGen Mathew et al. (2021) Det.
[B] Toxic, Benign
GPT-3 274,186
NACL Masud et al. (2022) Det.
[M.C.] Hate Intensity [M.L.] Hate Spans
Mixed 4,423
Ctr.
[F.T.] Hate Speech Normalization
CONAN Chung et al. (2019) Det.
[M.C.] Hate Types [M.C.] Hate Sub-Topic
Synthetic 14,988
Ctr.
[F.T.] CN Generation
Multitarget CONAN Fanton et al. (2021) Det.
[M.C.] Hate Targets
GPT-2 5,003
Ctr.
[F.T.] CN Generation
Counter Narratives Das et al. (2023) Ctr.
[F.T.] CN Generation
YouTube 9,119
Img MMHS150K Gomez et al. (2020) Det.
[B] Hateful
Twitter 150,000
FHM Kiela et al. (2020) Det.
[B] Hateful
Synthetic 10,000
Finegrained FHM Mathias et al. (2021) Det.
[B] Hateful [M.L.M.C] Protected Category [M.L.M.C] Protected Attacks
Synthetic 10,000
Misogynous Meme Det.
[B] Misogynistic [B] Aggressive
Mixed 800
Gasparini et al. (2022)
[B] Ironic
MAMI Fersini et al. (2022) Det.
[B] Misogyny [M.L.M.C.] Misogynous, Shaming, Stereotype, Objectification, Violence
Mixed 10,000
UA-RU Conflict Thapa et al. (2022) Det.
[B] Hateful
Twitter 5,680
CrisisHateMM Bhandari et al. (2023) Det.
[B] Hateful [B] Directed [M.C.] Hate Targets
Mixed 4,723
HatReD Hee et al. (2023) Expl.
[F.T] Explanations
Synthetic 3,228
Video HateMM Das et al. (2023) Det.
[B] Hateful [M.C.] Hate Targets
Mixed 1,083
Bangla Hate Videos Junaid et al. (2021) Det.
[B] Hateful
YouTube 300
Audio DeToxy Ghosh et al. (2021) Det.
[B] Hateful / Non-hateful
Mixed 2M
MuTox Costa-jussà et al. (2024) Det.
[B] Hateful / Non-hateful
Mixed 116,000
In this section, we review state-of-the-art methodologies that have made significant contributions to key areas of HS research: HS detection, HS explanation, HS debiasing, and counter speech. We emphasise research studies incorporating LLMs or LMMs, aiming to garner comprehensive insights into emerging trends within this domain. This allows us to understand how these advanced models enhance our understanding and management of HS in various forms.

### v4

@@ Abstract
Moderating hate speech (HS) in the evolving online landscape is a complex challenge, compounded by the multimodal nature of digital content. This survey examines recent advancements in HS moderation, focusing on the burgeoning role of large language models (LLMs) and large multimodal models (LMMs) in detecting, explaining, debiasing, and countering HS. We begin with a comprehensive analysis of current literature, uncovering how text, images, and audio interact to spread HS. The combination of these modalities adds complexity and subtlety to HS dissemination. We also identified research gaps, particularly in underrepresented languages and cultures, and highlight the need for solutions in low-resource settings. The survey concludes with future research directions, including novel AI methodologies, ethical AI governance, and the development of context-aware systems. This overview aims to inspire further research and foster collaboration towards responsible and human-centric approaches to HS moderation in the digital age.111 WARNING: This paper contains offensive examples.
Recent Advances in Online Hate Speech Moderation:
Multimodality and the Role of Large Models
Ming Shan Hee1thanks: These authors contributed equally to this work., Shivam Sharma2,*, Rui Cao3, Palash Nandi2, Preslav Nakov4, Tanmoy Chakraborty2, Roy Ka-Wei Lee1, 1SUTD, 2IIT Delhi, 3SMU, 4MBZUAI 1{mingshan_hee@mymail., roy_lee@}sutd.edu.sg 2{shivam.sharma, palash.nandi, tanchak}@ee.iitd.ac.in 3ruicao.2020@phdcs.smu.edu.sg 4preslav.nakov@mbzuai.ac.ae
@@ 1 Introduction
In the era of rapid information exchange and digital connectivity, the rise of hate speech (HS) presents a significant challenge with profound implications for global societies. HS, which is any communication demeaning a person or a group based on social or ethnic characteristics, undermines social harmony and individual safety, both online and offline Lupu et al. (2023). The recent Israel–Hamas conflict has notably escalated both anti-Muslim and anti-Semitic sentiments worldwide, evidenced by the trending of hashtags such as #HitlerWasRight and #DeathToMuslim on the social media platform X. [^2] https://www.nytimes.com/2023/11/15/technology/hate-speech-israel-gaza-internet.html Moreover, the Council on American–Islamic Relations reported receiving 774 help requests and bias reports from Muslims in the USA within a 16-day period. [^3] https://www.cair.com/press_releases/cair-reports-sharp-increase-in-complaints-reported-bias-incidents-since-107/ While digital interconnectivity facilitates swift information sharing, it simultaneously amplifies the spread and the impact of HS, transcending geographical boundaries.
@@
Technological advancements have transformed the expression of HS, leading to its manifestation in various novel forms. Traditionally, HS was predominantly text-based, found in written materials Rini et al. (2020), or verbalized in posts, broadcasts, and public speeches Nielsen (2002). The digital era has ushered in more complex and subtle variants of HS, engaging multiple sensory modalities. A notable instance is vision-language HS, which fuses visual elements with text, commonly disseminated through captioned images and memes Uyheng et al. (2020); Kiela et al. (2020). Video-based HS, another emerging form, amalgamates text, visuals, and audio, creating a multi-faceted and potentially more influential mode of communication Das et al. (2023). Figure 1 shows various HS forms targeting immigrants, underscoring animosity towards individuals of diverse nationalities. The text-based approach overtly projects hostile attitudes towards them in the host country. In vision-language HS, visual (e.g., a person preparing to shoot) and textual elements (e.g., sighting an illegal immigrant mowing the lawn) jointly convey antagonism. The figure also includes a music parody, integrating derogatory visuals with discriminatory audio lyrics, to showcase contempt for immigrants.
While existing research surveys Rini et al. (2020); Chhabra and Vishwakarma (2023); Subramanian et al. (2023) have largely focused on text-based HS, they often overlook the complexity of multimodal content. Our survey addresses this gap by offering a comprehensive analysis of HS across various digital platforms, including text, visual, auditory, and combined multimodal expressions. We explore the distinct ways HS manifests in these formats, providing insights into their characteristics and moderation challenges. Additionally, we emphasize the critical role of large language models (LLMs) and large multimodal models (LMMs) in moderating HS, given their ability to process and interpret diverse data types. This survey critically evaluates existing solutions, identifies areas for improvement, and advocates for a shift towards multimodal approaches in HS moderation.
@@
Paper Collection. We systematically examined research pertaining to the moderation of various types of hate speech, encompassing text, images, videos, and audio. Our search involved keywords such as ’hate speech’, ’multimodal hate speech’, ’hateful memes’, and similar terms, across scholarly platforms like Google Scholar, DBLP, IEEE Xplore, and ACM Digital Library. Among related research, we further selected state-of-the-art studies, with a particular interest in those using LLMs and LMMs. Due to the need for a manageable scope, we excluded works that did not leverage LLMs or LMMs, or focused narrowly on regional or multilingual aspects without broader relevance. This decision is not a reflection on the quality or importance of these works but rather a necessity to maintain a focused and coherent survey.
Mod. Dataset Task Labels Source # Records
Text WZ-LS Waseem and Hovy (2016) Det. [M.C.] Sexism, Racism, Neither Twitter 16,914
GHC Kennedy et al. (2018) Det. [M.C.] VO, HD, CV [B] Implicitness [M.C.] Hate Targets Forums 27,665
Stormfront de Gibert et al. (2018) Det. [B] Hateful StormFront 9,916
DT Davidson et al. (2017) Det. [M.C.] Hateful, Offensive, Neither Twitter 24,802
Founta Founta et al. (2018) Det. [M.C.] Offensive, Abusive, Hateful Speech, Aggressive, Cyberbullying, Spam, Normal Twitter 80,000
DynaHate Vidgen et al. (2021) Det. [B] Hateful, [M.C.] Hate Targets, [M.C] Animosity, Derogation, Dehumanization, Threatening, Support H-M Adv 41,134
SBIC Sap et al. (2020) Det. [B] Offensive [M.C] Hate Targets [B] Intent [B] Lewd [B] Group [B] In-Group Mixed 44,671
Expl. [F.T.] Implied Statement
IHC ElSherief et al. (2021) Det. [M.C.] Implicit, Explicit, Non-Hate [M.C.] Grievance, Incitement, Inferiority, Irony, Stereotypical, Threatening, Others Twitter 22,584
Expl. [F.T.] Implied Statement
HateXplain Mathew et al. (2021) Det. [M.C] Hate, Offensive, Normal [M.C.] Hate Targets Mixed 20,148
Expl. [M.L] Text Rationales/Snippets
NACL Masud et al. (2022) Det. [M.C.] Hate Intensity [M.L.] Hate Spans Mixed 4,423
Ctr. [F.T.] Hate Speech Normalization
CONAN Chung et al. (2019) Det. [M.C.] Hate Types [M.C.] Hate Sub-Topic Synthetic 14,988
Ctr. [F.T.] CN Generation
Multitarget CONAN Fanton et al. (2021) Det. [M.C.] Hate Targets GPT-2 5,003
Ctr. [F.T.] CN Generation
Counter Narratives Das et al. (2023) Ctr. [F.T.] CN Generation YouTube 9,119
Img MMHS150K Gomez et al. (2020) Det. [B] Hateful Twitter 150,000
FHM Kiela et al. (2020) Det. [B] Hateful Synthetic 10,000
Finegrained FHM Mathias et al. (2021) Det. [B] Hateful [M.L.M.C] Protected Category [M.L.M.C] Protected Attacks Synthetic 10,000
Misogynous Meme Det. [B] Misogynistic [B] Aggressive Mixed 800
Gasparini et al. (2022) [B] Ironic
MAMI Fersini et al. (2022) Det. [B] Misogyny [M.L.M.C.] Misogynous, Shaming, Stereotype, Objectification, Violence Mixed 10,000
UA-RU Conflict Thapa et al. (2022) Det. [B] Hateful Twitter 5,680
CrisisHateMM Bhandari et al. (2023) Det. [B] Hateful [B] Directed [M.C.] Hate Targets Mixed 4,723
RUHate-MM Thapa et al. (2024) Det. [B] Hateful [M.C] Hate Targets Twitter 20,675
HatReD Hee et al. (2023) Expl. [F.T] Explanations Synthetic 3,228
Video Bangla Hate Videos Junaid et al. (2021) Det. [B] Hateful YouTube 300
HateMM Das et al. (2023) Det. [B] Hateful [M.C.] Hate Targets Mixed 1,083
MultiHateClip Wang et al. (2024) Det. [M.C] Hateful, Offensive, Normal YouTube & Bilibili 2,000
Audio DeToxy Ghosh et al. (2021) Det. [B] Hateful Mixed 2M
MuTox Costa-jussà et al. (2024) Det. [B] Hateful Mixed 116,000
@@ 2 Hate Speech
Figure 2: Typology of HS based on modalities and tasks. The dark blue boxes are mature areas with multiple studies; light grey boxes are ongoing research, and hatched boxes are unexplored topics.
HS takes various forms — written text, images, spoken words, and multimedia content — each posing risks of violence, animosity, or prejudice against specific groups. This section reviews existing literature on HS, categorizing it into text-based, image-based, video-based, and audio-based types. For each HS form, we provide a detailed categorization across four tasks: detection, explanation, debiasing, and counter-speech. Detection identifies hateful content, forming the basis for further actions. Explanation promotes transparency by clarifying why content is flagged, building trust in automated systems. Debiasing is essential to refine detection systems, ensuring fairness and reducing bias. Counter-speech involves taking proactive steps to mitigate the impact of hate speech, fostering healthier online dialogue. Although these tasks address different aspects, they collectively form the foundation of an effective content moderation strategy, highlighting both the interconnectedness of HS forms and the research gaps in advancing multimodal HS moderation. Figure 2 illustrates the range of online HS forms. Additionally, Table 1 lists publicly accessible HS datasets in different modalities, providing researchers with essential resources for HS moderation.
@@ 2.1 Text-based Hate Speech
Text-based HS encompasses written or typed expressions manifested across online platforms, such as social media posts Waseem and Hovy (2016); Founta et al. (2018). Recent studies explored diverse aspects of hate and derogatory language, focusing on implicit HS Sap et al. (2020), targeted groups Kennedy et al. (2018); Yoder et al. (2022), and types of attacks ElSherief et al. (2021). As HS detection models improve, it becomes imperative to understand and explain their decision-making processes, mitigating unintended bias Garg et al. (2023). Additionally, some research shifted towards proactive strategies, including countering HS Masud et al. (2022).
The detection of text-based HS poses numerous challenges. Detecting hate speech (HS) in a single statement often requires understanding dark humor and cultural nuances Hee et al. (2024). HS can express underlying intent through sarcasm, irony, or cultural references, which may not be immediately apparent. Linguistic variations, such as slang, dialects, and unconventional language use, further complicate the task. The challenge intensifies when considering the broader context of an utterance Nagar et al. (2023); Yu et al. (2022a), as statements that seem neutral in isolation may reveal hateful intent when viewed within a conversation. Conversely, what appears offensive might be harmless in context. Therefore, context-aware models are essential for accurately identifying HS by analyzing both individual statements and their surrounding situational context. Expanding the analysis to converssations, such as Reddit threads or WhatsApp chat, adds additional layers of complexity Naseem et al. (2019). The intent behind a single message can shift based on prior exchanges and the overall tone of the conversation. Furthermore, user-specific features may be important for HS detection Qian et al. (2018). Data such as a user’s posting history, profile, and behavior provide valuable context for identifying hate speech, though using such data raises ethical concerns, particularly regarding privacy.
@@ 2.2 Image-based Hate Speech
Image-based HS utilizes visual elements, such as photographs, cartoons, and illustrations, to propagate hate or discrimination against specific groups. A common manifestation of this HS form is memes, which typically consist of images combined with short overlaid text. Although memes often serve humorous or satirical purposes, they are increasingly used to spread hateful content online Kiela et al. (2020). Recent studies have developed datasets for identifying HS Gomez et al. (2020), specific targets Mathias et al. (2021) and types of attacks Fersini et al. (2022) within these memes. Beyond detection, new approaches analyze and mitigate bias in image-based HS detection models Hee et al. (2023); Lin et al. (2023). Additionally, new methodologies are emerging to counteract HS transmitted through memes Van and Wu (2023).
Image-based HS presents new challenges due to the subtlety of offensive messages concealed within multiple modalities. Images, often embedding symbols, memes, or culturally specific visual cues, require deep cultural and contextual understanding for accurate interpretation. The visual elements and text can subtly imply meanings not immediately evident Kiela et al. (2020). For example, Figure 1 depicts a man with a gun and text suggesting hostility towards immigrants. Differentiating humour from hate in memes is particularly challenging, influenced by varying cultural, societal, and personal perspectives Schmid (2023).
@@ 2.3 Video-based Hate Speech
Video-based HS presents a complex challenge, comprising a blend of visuals, audio tracks, and/or textual elements. This form of HS ranges from professionally produced propaganda to amateur videos on social media platforms like YouTube and TikTok Das et al. (2023); Wang et al. (2024). The engaging nature of video content and its easy dissemination across digital networks significantly heighten its potential for harm. Echoing the concerns of image-based HS, video-based HS also contributes to the normalization of hateful ideologies and can profoundly influence public opinion. Contemporary research primarily focuses on identifying video-based HS and categorizing its various subtypes Wu and Bhandary (2020). Nonetheless, the amount of research on video-based HS is less developed than text-based and image-based HS, particularly in areas such as analyzing and mitigating model bias, elucidating decision-making processes, and devising counterstrategies. These gaps, likely stemming from the rapid pace of technological advancements and evolving digital trends, underscore the need for further research to promote a more harmonious online environment.
Detecting hate speech (HS) in videos is challenging and resource-intensive because it requires understanding various elements, including text, images, and audio, both independently and in combination. Each component can independently contain hateful content, further complicating the detection process. The duration of videos further exacerbates this challenge, as longer content necessitates more extensive review and analysis, with potential shifts in context over time. Moreover, subtle visual cues and sophisticated editing techniques can be employed to discreetly embed hate messages, making their detection by automated tools particularly challenging. Additionally, video content analysis requires considerable computational resources and time, posing a substantial challenge for organizations to detect and address HS in video formats.
@@ 2.4 Audio-based Hate Speech
Audio-based HS entails the analysis of sound waves to discern elements such as pitch, intonation, and the contextual meaning of spoken words. This form of HS can originate from a variety of audio channels, including real-time conversations, podcasts, and other forms of audio media. The methodologies for addressing audio-based HS are diverse, targeting different facets of the issue. For instance, Barakat et al. (2012) employed a straightforward keyword-based approach to identify segments of HS, while Wazir et al. (2020) engaged in a detailed classification of offensive categories in audio-based HS, showcasing a nuanced method of understanding and categorizing this form of HS. This research area is still in its developmental stages, partly due to the scarcity of dataset. Nonetheless, recognizing the variety and significance of the approaches and techniques employed in this field is imperative. This recognition not only sheds light on the current state of research but also illuminates potential avenues for future exploration.
@@ 3 Methodology
This section reviews the state-of-the-art methodologies that have significantly contributed to primary areas of HS research, particularly those involving large models. First, we discuss the recent capabilities of large models (Section 3.1). Subsequently, we explore studies in four important HS areas: detection (Section 3.2), explanation (Section 3.3), debiasing (Section 3.4), and counter-speech (Section 3.5), focusing on works using large models. This review highlights the emerging trends, providing insights into how large models can be used to understand and address HS in its various forms.